OpenStudy (anonymous):

What are the solutions of the equation below: z²-6z-27=0 Please Answer!! =)

OpenStudy (anonymous):

\[z^2 - 6z -27 = (z-9)(z+3)\] z= 9 or z = -3

OpenStudy (anonymous):

z(z-9)+3(z-9) (z+3)(z-9) z=-3 or z=9